# Branch length statistics in phylogenetic trees under constant-rate birth-death dynamics

### v3

@@ V Branch length distribution: interior branches
As a next step, we derive the distribution of internal branch lengths, i.e., all branches that are not pendant. (We exclude the branch between the start of the birth-death process to the first bifurcation.) To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
Two cases must be considered for the other clade (called $B$) emerging at $\tau_{s}$. With probability $p_{1}(\tau_{s})$ this clade becomes a pendant branch. In this case, a factor of two arises due to symmetry (either clade can be called $A$). An alternative scenario is the clade having more than one extant descendant. This happens with probability $p_{>1}(\tau_{s})=1-p_{0}(\tau_{s})-p_{1}(\tau_{s})$. In this case, the clade also has an interior branch starting at $\tau_{s}$ contributing to the branch length distribution. A factor of two accounts for this.
Putting these steps together, the relative probability of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalization factor found to be one) is given by
$P_{\text{int}}(\tau_{s},\tau_{e})$ $=\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}\left[2p_{1}(\tau_{s})+2p_{>1}(\tau_{s})\right]$
$=2\lambda^{2}\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\left(1-p_{0}(\tau_{e})\right)^{2}\left(1-p_{0}(\tau_{s})\right)$ (13)
$=2\lambda^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}-\mu)^{3}}\ .$
@@
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 00, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13). Again, this factor decays exponentially towards one as the population age $T$ increases, justifying the asymptotic results also as an approximation for populations of finite age with $(\lambda-\mu)T\gg 1$.
@@ VI Finite sampling
Usually only a subset of individuals sampled from the population is available to construct the phylogenetic tree from and this sampling affects the statistics of branch lengths. To model the process of sampling from the extant population we follow the standard approach and choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalized to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
@@
[…]
@@
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. In particular, we obtain
@@
A normalizing factor of $\rho^{-1}$ applies to all three densities and is included in the expressions above. The result for pendant branches coincides with the result of Stadler and Steel Stadler and Steel (2012) at fixed number of samples. However, their result was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (18) also holds outside these limits.
@@
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((18) and (20) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the forth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (21) (red dashed line) as $\rho$ becomes small. (B) A toy example of how sampling affects the length of pendant branches, see text.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to a particular set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (18)-(20), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument used in Stadler (2009, 2013a)). Therefore, inference of both growth parameters and the sampling probability from the distribution of branch lengths is impossible. However, an upper bound on the sampling probability can be determined: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of possible parameter sets.
@@
For any finite $\tau_{e}$, this distribution approaches zero with decreasing sampling probability $\rho$. This means that the ends of interior branches lie asymptotically far away from the time of observation when $\rho$ approaches zero. Setting equation (22) equal to $1/2$ and solving for $\tau_{e}$ yields the median interior branch end. In the regime $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}\ll 1$, the median is given by $\ln\left((\lambda-\mu)/\rho\right)/(\lambda-\mu)$, reproducing the dependence on the sampling probability found in the toy model.
@@ VII Branch lengths conditioned on the population size at the time of observation
We first illustrate the method with a toy problem: We calculate the distribution of the length of pendant branches when the number of individuals at the time of observation is $N=2$. A key tool is the probability that a birth-death process with birth rate $\lambda$ and death rate $\mu$, starting with $n$ individuals at a given time has $N$ individuals a time interval $t$ later, denoted $p_{N,n}^{\lambda,\mu}(t)$.
Figure 4: Pendant branch formation for populations of size $N=2$: The red branch marks the pendant branch arising from a birth event at time $\tau$ in the past. All other $n-1$ individuals that were alive at time $\tau$ need to die out before the population is observed, while the second descendant from the birth event needs to survive with clade size one. The arrows on the right illustrate the conditional probabilities of having $N$ extant descendants given that $n$ individuals at time $\tau$ before observation (forwards process) and the probability that given $N$ observed individuals there were $n$ individuals at time $\tau$ before observation (backwards process).
First we go backwards in time from the time of observation to the time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation has been characterized by Ignatieva, Hein and Jenkins Ignatieva et al. (2020). It is described by that of a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{n,N}^{\mu,\lambda}(\tau)$. The number of individuals $n$ will eventually be summed over.
@@
$\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)/p_{2,n}^{\lambda,\mu}(\tau)\ .$ (23)
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other $n-1$ individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N=2$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N=2$ individuals at the time of observation. For an illustration, see Figure 4.
@@
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{n,2}^{\mu,\lambda}(\tau)}{p_{2,n}^{\lambda,\mu}(\tau)}\ .$ (24)
In Appendix .1 we show how this expression evaluates to $2\lambda\frac{p_{1}^{2}}{1-p_{0}}=2\lambda\frac{(\lambda-\mu)^{3}\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}}{\left(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau}-\mu\right)^{3}}$ and a normalisation constant of one. We find the same result for a final population size $N\geq 2$, see Appendix .1. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $\tau$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Running time forwards again while conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
The resulting distribution of interior branch lengths under scenario ii) depends on $N$. However, the dependency quickly vanishes with $\left(\frac{\lambda}{\mu}p_{0}(\tau_{s})\right)^{N-1}$ when increasing the population size. The asymptotic distribution for large $N$ coincides with the interior branch length distribution under scenario i) in the limit of large times, see Appendix .2.

### v4

@@ V Branch length distribution: interior branches
As a next step, we derive the distribution of internal branch lengths, i.e., all branches that are not pendant (nor the root branch leading from the start of the birth-death process to the first bifurcation). To this end, we calculate the joint probability of an interior branch starting at time $\tau_{s}$ and ending at $\tau_{e}<\tau_{s}$. As before, the probability per unit time of a birth event occurring at time $\tau_{s}$ is proportional to $\lambda\operatorname{\mathrm{e}}^{-(1-q)\tau_{s}}$. Focusing for now on only one of the two clades that started at $\tau_{s}$ (called $A$), the probability of it running until $\tau_{e}$ with only one extant descendant at the time of observation is $p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})$. The probability per small time interval of the branch ending at $\tau_{e}$ is equal to the probability per small time interval of a birth event at that time where none of the two emerging clades dies out. This is given by $\lambda(1-p_{0}(\tau_{e}))^{2}$.
The second clade (called $B$) emerging at $\tau_{s}$ must not die out, which happens with probability $1-p_{0}(\tau)$. The probability that both branches emerging at $\tau_{s}$ end at a small interval around $\tau_{e}$ vanishes quadratically with the time interval and can be neglected. Either clade can be called $A$, which leads to an overall factor of two.
Putting these elements together, the probability density function of an interior branch running from $\tau_{s}$ to $\tau_{e}$ (up to a normalization factor found to be one) is given by
$P_{\text{int}}(\tau_{s},\tau_{e})$ $=\lambda\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{s}}p_{1}^{(\tau_{e})}(\tau_{s}-\tau_{e})\lambda\left(1-p_{0}(\tau_{e})\right)^{2}2\left(1-p_{0}(\tau_{s})\right)$
$=2\lambda^{2}(\lambda-\mu)^{3}\frac{\operatorname{\mathrm{e}}^{(\lambda-\mu)(\tau_{s}+\tau_{e})}}{(\lambda\operatorname{\mathrm{e}}^{(\lambda-\mu)\tau_{s}}-\mu)^{3}}\ .$ (13)
@@
So far, we have again only considered the asymptotic case of $T\rightarrow\infty$. For finite $T$, the probability of $\tau_{s}>T$ is 0, resulting in a normalization factor of $(\lambda-\mu\operatorname{\mathrm{e}}^{-(\lambda-\mu)T})^{2}/(a(1-\operatorname{\mathrm{e}}^{-(\lambda-\mu)T}))^{2}$ for (13). Again, this factor decays exponentially towards one as the population age $T$ increases, justifying the asymptotic results also as an approximation for populations of finite age with $(\lambda-\mu)T\gg 1$.
@@ VI Finite sampling
Usually, only a subset of individuals sampled from the population is available to construct the phylogenetic tree from, and this sampling affects the statistics of branch lengths. To model the process of sampling from the extant population we follow the standard approach and choose extant individuals at the time of observation independently and with equal probability $\rho$ Yang and Rannala (1997); Stadler (2010); Stadler and Steel (2012). The preceding results on branch length distributions can be easily generalized to a sampling probability $\rho<1$ using the probabilities that a clade of a particular age has no (or one) member which is sampled at a particular time.
@@
These results can be derived by summing over the probabilities of clade sizes $n$ at time $\tau$ (equation (4)) weighted with $(1-\rho)^{n}$ and $n\rho(1-\rho)^{n-1}$, respectively.
Based on these probabilities, we can repeat the derivations from Section II to compute the probability $p_{1}^{(\rho)(\tau_{e})}(\tau)$ that a clade starting with a single individual $\tau_{e}+\tau$ before the time of observation has one descendant alive at time $\tau$ (with an open future fate) and none of the additional lineages arising between $\tau_{e}+\tau$ and $\tau$ leave are sampled at time 0. Replacing $p_{0}(\tau)$ with (15) in master equation (6) yields
@@
To derive the probability density function of branch lengths under sampling, these probabilities can be used directly in the densities of pendant and interior branch lengths at full sampling ($\rho=1$), (10)-(13), which depend only on $p_{0}$, $p_{1}$, and the relative population sizes at different times. Replacing $p_{0}(\tau)$, $p_{1}(\tau)$ and $p_{1}^{(\rho)}(\tau)$ with their respective counterparts (15)-(17) thus fully accounts for finite sampling, as the population size in the past is not affected by sampling of the extant population. In this way, we obtain
@@
A normalizing factor equalling $\rho^{-1}$ in the limit of large $T$ applies to all three densities and is included in the expressions above. The result for pendant branches coincides with the result of Stadler and Steel Stadler and Steel (2012) at a fixed number of samples. However, their result was restricted to a regime of high death rate or high sampling probability $1\geq\mu/\lambda\geq 1-\rho$. The density function (18) also holds outside these limits.
@@
Figure 3: Pendant and interior branches at different sampling probabilities $\rho$. (A) We compare the analytical results for the probability density of the length of pendant and interior branches under sampling ((18) and (20) respectively, solid blue lines) to numerical simulations. For the latter, we show the distribution of branches from 100 trees. For every tree, we simulated the growth of a population with $\lambda=1$ and $\mu=0.75$ over time $T=40$, restarting when the population died out or the final population size was not within $2\times 10^{4}$ and $3\times 10^{5}$ (grey bars). The first row of plots shows the branch length distribution of the full trees. For the next row of plots, individuals were sampled with probability $\rho=0.1$ and the corresponding trees were reconstructed. For the third row, each of the previous samples was sampled again with probability $0.1$, corresponding to an overall sampling probability of $\rho=0.01$. The same was repeated for the fourth row of plots, yielding $\rho=0.001$. Every tree contributes equally to the empirical distributions depicted. The first column shows the distributions of pendant branch lengths, which shift to longer branches as $\rho$ decreases. The second column of plots shows the distribution of interior branches, quickly reaching the asymptotic distribution (21) (red dashed line) as $\rho$ becomes small. (B) A toy example of how sampling affects the length of pendant branches, see text.
@@
For any finite $\tau_{e}$, this distribution approaches zero with decreasing sampling probability $\rho$. This means that the ends of interior branches lie asymptotically far away from the time of observation when $\rho$ approaches zero. Setting equation (22) equal to $1/2$ and solving for $\tau_{e}$ yields the median interior branch end. In the regime $\operatorname{\mathrm{e}}^{-(\lambda-\mu)\tau_{e}}\ll 1$, the median is given by $\ln\left((\lambda-\mu)/\rho\right)/(\lambda-\mu)$, reproducing the logarithmic dependence on the sampling probability found in the toy model.
Stadler Stadler (2009) found that the probability distribution of bifurcation times is not unique to a particular set of parameters, but that a birth-death process with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho$ yields the same distribution as a process with parameters $\rho^{\prime}$, $\lambda^{\prime}=\rho\lambda/\rho^{\prime}$, $\mu^{\prime}=\mu-\lambda(1-\rho/\rho^{\prime})$. We find the same for the distributions of branch lengths (18)-(20), as $\lambda^{\prime}-\mu^{\prime}=\lambda-\mu$, $\rho^{\prime}\lambda^{\prime}=\rho\lambda$ and $\lambda^{\prime}(1-\rho^{\prime})-\mu^{\prime}=\lambda(1-\rho)-\mu$ (which is the same argument used in Stadler (2009, 2013a)). Therefore, inference of both growth parameters and the sampling probability from the distribution of branch lengths is impossible. However, an upper bound on the sampling probability can be determined: Given a branch distribution generated with (unknown) parameters $(\lambda,\mu,\rho)$, the parameter space of all sets of $(\lambda^{\prime},\mu^{\prime},\rho^{\prime})$ which lead to the same distribution is restricted to $\mu^{\prime}\geq 0$ and $0<\rho\leq 1$. With the transformation rules above, it is easy to see that these inequalities are only fulfilled if $0<\rho^{\prime}\leq\min(\rho\lambda/(\lambda-\mu),1)$. This means that for branch distributions generated with birth rate $\lambda$, death rate $\mu$ and sampling probability $\rho<1-\mu/\lambda$, there exists a maximal sampling probability $\rho^{\prime}<1$ for the family of possible parameter sets. This agrees qualitatively with the behavior of pendant branches in Figure 3A: At full sampling, the maximum of the pendant branches is always found at $\tau=0$. With decreasing sampling probability the time $\tau$ of the maximum eventually becomes non-zero and continues to increase monotonically when the sampling probability is decreased further. For a tree with a pendant branch distribution with a maximum at $\tau>0$, no set of parameters $\lambda,\mu$ can be found that together with $\rho=1$ would lead to the observed distribution.
@@ VII Branch lengths conditioned on the population size at the time of observation
We start by illustrating the method with a toy problem: We calculate the distribution of the length of pendant branches when the number of individuals at the time of observation is $N=2$. A key tool is the probability that a birth-death process with birth rate $\lambda$ and death rate $\mu$, starting with $n$ individuals has $N$ individuals after evolving for time $t$, denoted $p_{(N,n)}^{(\lambda,\mu)}(t)$.
Figure 4: Pendant branch formation for populations of size $N=2$: The red branch marks the pendant branch arising from a birth event at time $\tau$ in the past. All other n1n-1 individuals that were alive at time $\tau$ need to die out before the population is observed, while the second descendant from the birth event needs to survive with clade size one. The arrows on the right illustrate the conditional probabilities of having $N$ extant descendants given that $n$ individuals at time $\tau$ before observation (forwards process) and the probability that given NN observed individuals there were $n$ individuals at time $\tau$ before observation (backwards process).
We start by considering the population at the time of observation (with fixed population size $N$) and ask about its population size at a previous time $\tau$. Conditioned on the number of individuals at the time of observation being $N$, the statistics of the number of individuals before the time of observation has been characterized by Ignatieva, Hein and Jenkins Ignatieva et al. (2020). It is described by that of a birth-death process with time running backwards and with the rates of birth and death exchanged, so the rate of birth is $\mu$ and the rate of death is $\lambda$ Ignatieva et al. (2020). The probability of having $n$ individuals at time $\tau$ before the time of observation is thus $p_{(n,N)}^{(\mu,\lambda)}(\tau)$. This number of individuals $n$ will eventually be summed over.
@@
$\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau){\frac{1}{p_{(2,n)}^{(\lambda,\mu)}(\tau)}}.$ (23)
$\lambda n$ gives the probability per small time interval that one of the $n$ individuals divides during this small interval, $p_{1}^{2}(\tau)$ is the probability that the two resulting offspring each have one extant offspring at the time of observation, leading to pendant branches of length $\tau$. $p_{0}^{n-1}(\tau)$ is the probability that all other n1n-1 individuals die out by the time of observation, leading to a population of size $N=2$. The denominator arises from the conditioning on $N=2$ individuals. What we have done is first asking about the statistics of the number of individuals at a certain time $\tau$ in the past, given $N=2$ individuals at some time of observation, and then asking about the probability that a tree with a particular feature (here a pendant branch that emerged at time $\tau$) arises from those forward-time trajectories ending in $N=2$ individuals at the time of observation. We thus first went ”backwards in time” from the time of observation to a time $\tau$ before that, and then ”forwards in time” to the time of observation. For an illustration, see Figure 4.
@@
$\sum_{n=1}^{\infty}\lambda np_{1}^{2}(\tau)p_{0}^{n-1}(\tau)\frac{p_{(n,2)}^{(\mu,\lambda)}(\tau)}{p_{(2,n)}^{(\lambda,\mu)}(\tau)}{=2\lambda p_{1}(\tau)^{2}\sum_{n=1}^{\infty}p_{0}(\tau)^{n-1}=2\lambda\frac{p_{1}(\tau)^{2}}{1-p_{0}(\tau)}}\ .$ (24)
Here, we have used that $\frac{p_{(n,2)}^{(\mu,\lambda)}(\tau)}{p_{(2,n)}^{(\lambda,\mu)}(\tau)}=\frac{N}{n}$, which we show in Appendix .1 (equation (26)). The normalizing constant turns out to be one.
We find the same result for a final population size $N\geq 2$, see Appendix .1. This reproduces the result by Stadler and collaborators on the distribution of pendant branches conditioned on the size $N$ of the population at the time of observation. Curiously, this distribution also coincides with the distribution of pendant branch lengths under scenario i) in the limit of large $T$ (equation (10)).
The same approach can also be used to calculate the distribution of interior branch lengths conditioned on the population size. Again we use the birth-death process with time first running backwards to compute the probability that at a time $\tau_{s}$ in the past there were $n$ individuals, one of which divided establishing a new branch. Then running time forwards again while conditioning on the final population size $N$, we calculate the probability that the new branch ends at $\tau_{e}$ in another birth event, with both descendants having extant offspring. The conditions that lead to a branch of a particular length are the same manner as in scenario i), the difference is the conditioning on the final population size in two directions of time. The calculation is given in Appendix .2.
The resulting distribution of interior branch lengths under scenario ii) depends on $N$. However, the dependency quickly vanishes with $\left(\frac{\lambda}{\mu}p_{0}(\tau_{s})\right)^{N-1}$ when increasing the population size. The asymptotic distribution for large $N$ coincides with the interior branch length distribution under scenario i) in the limit of large times $T$, see Appendix .2.